OpenStudy (anonymous):

Using the letters of the word YOUNG, tell how many different 5-letter combinations are possible if: the first letter must be Y the vowels and consonants alternate, beginning with a consonant (Y is a consonant here and Y does not have to be first).

OpenStudy (anonymous):

i think 80, but could be wrong this time. since i have no clue what the vocabs mean in the third sentance

ganeshie8 (ganeshie8):

two questions here. lets answer one question at a time

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

1) the first letter must be Y you fix the first space as Y. remaining 4 spaces can be filled in 4*3*2*1 ways = 4!

ganeshie8 (ganeshie8):

2) must begin with consonant. and vowels and consonants alternate (vowels, consonants must not be together)

ganeshie8 (ganeshie8):

3 consonants , 2 vowels you can choose first letter from 3 consonants in : 3 ways after that, you can choose second letter from 2 VOWELS in : 2 ways after that, you can choose third letter from remaining 2 consonants in : 2 ways after that, you can choose fourth letter from remaining 1 VOWELS in : 1 ways finally, you can choose third letter from remaining 1 consonants in : 1 ways so total possible ways = 3*2*2*1*1 = 12

OpenStudy (anonymous):

so the answeres r 4 and 12

ganeshie8 (ganeshie8):

for first question, its 4! its not 4

ganeshie8 (ganeshie8):

did u understand how we worked it ?

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

good to hear that :P

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

tag me if u have more problems of this kind :)

OpenStudy (anonymous):

lol ok thanks!